# The possible adelic indices for elliptic curves admitting a rational cyclic isogeny

### v1

@@ 8.2. Rational Isogenies of Degree 22
24.36.0.cg.1 48.144.9.bep.1, 48.144.9.bfa.1, 48.144.9.bjz.1, 48.144.9.bks.1, 24.144.9.ejb.1, 48.144.7.bfg.1, 48.144.7.bft.1, 48.144.11.sm.1, 24.144.9.eid.1, 24.288.17.gin.1, 24.288.17.gil.1, 24.288.17.gim.1, 24.288.17.gim.2, 48.288.21.gvg.1, 48.288.21.gvf.1, 48.288.21.gvg.2, 48.288.21.gvh.1, 48.144.11.sr.1, 24.144.9.dga.1, 48.144.7.zv.1, 24.144.9.dgq.1, 48.144.7.ze.1
24.36.0.cj.1 48.288.15.cyo.1, 48.288.15.cva.1, 48.288.15.cyy.1, 48.288.15.cyv.1, 48.288.15.cvz.1, 48.288.15.cwc.1, 48.288.15.cyp.1, 48.288.15.cvb.1
24.36.1.gl.1 48.288.21.ezh.1, 48.288.17.fay.1, 48.288.19.hit.1, 48.288.19.hiq.1, 48.288.15.ceb.1, 48.288.15.cdy.1, 48.288.17.faz.1, 48.288.21.ezi.1
48.72.0.c.1 48.144.5.ex.1, 48.144.5.eu.1, 48.144.5.et.2, 48.144.5.eq.2, 48.144.3.z.1, 48.144.5.eh.2, 48.144.3.r.1, 48.144.11.ys.1, 48.144.5.em.1, 48.144.5.fb.2, 48.144.5.ei.2, 48.144.3.bb.1, 48.144.5.el.2, 48.144.11.yq.1, 48.144.5.ey.2, 48.144.5.ep.1, 48.144.5.eo.2, 48.144.5.ej.1, 48.144.5.ez.1, 48.144.5.en.2, 48.144.7.bas.1, 48.144.5.fa.1, 48.144.5.ek.1, 48.144.7.bgq.2, 48.144.5.es.1, 48.144.5.ew.2, 48.144.5.ef.1, 48.144.5.ee.2, 48.144.5.ev.2, 48.144.5.er.1, 48.144.3.t.2, 48.144.5.eg.1
48.72.0.c.2 48.144.5.ee.1, 48.144.7.bas.2, 48.144.3.z.2, 48.144.7.bgq.1, 48.144.3.r.2, 48.144.5.ek.2, 48.144.5.eu.2, 48.144.5.ei.1, 48.144.5.eq.1, 48.144.5.eo.1, 48.144.5.es.2, 48.144.5.em.2, 48.144.5.et.1, 48.144.5.ev.1, 48.144.5.ej.2, 48.144.5.ex.2, 48.144.5.ep.2, 48.144.5.er.2, 48.144.5.en.1, 48.144.3.bb.2, 48.144.5.el.1, 48.144.11.ys.2, 48.144.5.eh.1, 48.144.5.ez.2, 48.144.5.ef.2, 48.144.5.fb.1, 48.144.5.ey.1, 48.144.5.eg.2,48.144.5.fa.2, 48.144.5.ew.1, 48.144.11.yq.2, 48.144.3.t.1
48.72.0.d.1 48.144.5.fz.1, 48.144.5.gb.1, 48.144.5.gj.1, 48.144.5.fw.1, 48.144.3.bf.1, 48.144.3.bd.2, 48.144.5.gy.1, 48.144.5.hb.1, 48.144.5.fi.1, 48.144.5.gw.1
48.72.0.d.2 48.144.3.bf.2, 48.144.5.gd.1, 48.144.5.ga.1, 48.144.5.fx.1, 48.144.5.gf.1, 48.144.5.gu.1, 48.144.5.gx.1, 48.144.5.ha.1, 48.144.5.fm.1, 48.144.3.bd.1
First, many of these curves cover a rank 0 elliptic curve which has no non-CM rational points. Table LABEL:tab:coverings shows, for each of these, the elliptic curve covered.
Table 6. Coverings of rank 0 elliptic curves
Higher Genus Curve Rank 0 Elliptic Curve Covered
48.144.9.bep.1 12.18.1.g.1
48.144.5.gb.1 48.144.5.fx.1 12.36.1.bx.1
48.144.5.gw.1 48.144.5.ha.1 12.36.1.bu.1
48.288.15.cyo.1 48.144.5.gk.1 48.144.5.fw.1 48.144.5.ga.1 12.36.1.bw.1
48.288.21.ezh.1 48.288.17.fay.1 48.288.17.faz.1 48.288.21.ezi.1 12.36.1.bn.1
48.288.15.cyv.1 48.144.5.gu.1 24.36.1.gi.1
48.144.7.zv.1 48.144.7.ze.1 48.144.3.t.1 48.144.3.t.2 24.72.1.ef.1
48.288.15.ceb.1 48.288.15.cdy.1 48.144.5.gd.1 48.144.5.fz.1 24.36.1.gb.1
48.144.3.z.2 24.36.1.gq.1
48.144.3.z.1 24.36.1.gn.1
48.144.5.gj.1 48.144.5.ge.1 48.144.5.gf.1 24.36.1.ge.1
48.144.5.eg.2 48.24.1.j.1
48.144.5.gy.1 48.24.1.n.2
Meanwhile, other curves cover rank 0 curves of genus 2, and we can employ Chabauty0. These are shown in Table LABEL:tab:coverings2. We next highlight the non-CM rational points found. All the others had only cusps or CM points. First, the curves 48.144.5.em.2, 48.144.5.es.2, 48.144.5.eq.2, and 48.144.5.eo.2 cover 24.72.2.hl.1 which has genus 2, rank 0. Using Chabauty0, we check that this curve has exactly four rational points, two of which are cusps and two correspond to non-CM $j$-invariant $17^{3}$. The adelic index of the $j$-invariant $17^{3}$ is 1152. Also, the curves 48.144.5.eq.1, 48.144.5.eo.1, 48.144.5.em.1, and 48.144.5.es.1 all cover 24.72.2.hl.2 which has genus 2, rank 0. It has exactly four rational points: two are cusps and two correspond to the non-CM $j$-invariant $257^{3}$ that has adelic index 1152.
Table 7. Coverings applicable for Chabauty0
Higher Genus Curve Genus 2 Curve Covered
48.144.5.ex.2 48.144.5.ej.2 48.144.5.el.2 48.144.5.ev.2 24.72.2.hn.2
48.144.5.fa.2 48.144.5.eg.1 24.72.2.hj.2
48.144.5.ev.1 48.144.5.el.1 48.144.5.ej.1 48.144.5.ex.1 24.72.2.hn.1
48.144.5.eq.2 48.144.5.eo.2 48.144.5.em.2 48.144.5.es.2 24.72.2.hl.1
48.144.5.eq.1 48.144.5.eo.1 48.144.5.em.1 48.144.5.es.1 24.72.2.hl.2
48.144.5.fa.1 24.72.2.hj.1
@@
The curve 48.144.3.bf.1 is a genus 3 rank 3 curve. Instead of computing its set of rational points explicitly, we check that it has commutator index 576 and compute its proper maximal subgroups that are fiber products of two respective factors, contain I-I, and have full determinant. There are 11 such groups whose labels are in the set {48.288.13.bhs.1, 48.288.13.bhr.1, 48.288.13.bib.1, 48.288.13.bhy.1, 48.288.21.hxr.1, 48.288.13.bft.1, 48.288.21.hxt.1, 48.288.13.bfu.1, 48.288.13.bga.1, 48.288.13.bgd.1, 48.288.21.hxl.2}. Each of these can be handled using methods already discussed. The curves 48.288.13.bhs.1, 48.288.13.bhr.1, 48.288.13.bib.1, and 48.288.13.bft.1 cover 48.144.5.gh.1, 48.144.5.ge.1, 48.144.5.gk.1, and 48.144.5.fw.1 respectively, whose rational points we have already established. The curves 48.288.13.bhy.1, 48.288.21.hxt.1, 48.288.13.bga.1, 48.288.13.bgd.1 and 48.288.21.hxl.2 have no $\mathbb{Q}_{5}$ points. Lastly, the curves 48.288.21.hxr.1 and 48.288.13.bfu.1 have no real points.
@@
The curve 48.144.3.bd.1 is a genus 3, rank 3 curve. Again, instead of computing its set of rational points explicitly, we check that it has commutator index 576 and compute its proper maximal subgroups that are fiber products of two respective factors that contain I-I and have full determinant. There are 11 such curves and their labels comprise the set {48.288.21.hxh.2, 48.288.13.bfj.1, 48.288.13.bfk.1, 48.288.13.bgs.1, 48.288.13.bgr.1, 48.288.13.bfa.1, 48.288.13.bgi.1, 48.288.13.bfd.1, 48.288.13.bgl.1, 48.288.21.hxf.1, 48.288.21.hxp.2}. We now discuss their rational point computations. The curves 48.288.21.hxh.2, 48.288.13.bfj.1, 48.288.13.bfk.1, 48.288.13.bgs.1, 48.288.13.bgr.1 and 48.288.21.hxp.2 have no $\mathbb{Q}_{5}$ points. The curve 48.288.13.bfa.1 has no $\mathbb{Q}_{7}$ points. The curves 48.288.13.bgi.1, 48.288.13.bgl.1 and 48.288.21.hxf.1 have no real points. The curve 48.288.13.bfd.1 covers 24.36.1.fr.1, which has rank 0, and has exactly two rational points. One is a cusp and one is CM.
Finally, we discuss the curves 48.72.3.a.1 (= 16.24.0.k.2 $\times\;K_{2}$) and 48.72.3.a.2 (= 16.24.0.k.1 $\times\;K_{2}$). Let us consider 48.72.3.a.1 first. We divide our analysis into three parts. We first consider the direct product of children of 16.24.0.k.2 with $K_{2}$. The children of 16.24.0.k.2 are given by 16.48.0.n.1, 16.48.1.bq.1, 16.48.0.a.1, 16.48.0.h.2, 16.48.0.m.1, and 16.48.1.bv.1. The labels of the direct product of these with $K_{2}$ are 48.144.8.ca.1,
@@
48.144.7.rm.1, 48.144.8.a.1, 48.144.8.bd.1, 48.144.8.bz.2, and 48.144.7.sg.1} respectively. Let us discuss their rational points. The curve 48.144.8.ca.1 covers 12.36.1.bt.1 which has rank 0. It has exactly 6 points and all are cusps or CM. The curve 48.144.7.rm.1 covers 24.72.2.jo.1 which is bielliptic and has rank 2. It has Weierstrass model
@@
We next consider the direct product of 16.24.0.k.2 with the proper subgroups of $K_{2}$ in this lattice. The only such subgroup is $K_{3}$. The direct product of 16.24.0.k.2 with $K_{3}$ has label 48.144.7.ik.1. This curve covers 12.36.1.bq.1 which we have already discussed: it has no non-CM rational points.
@@
We now discuss their rational points. The curve 48.144.7.hu.1 covers 12.36.1.bn.1. By Table LABEL:tab:coverings, it has no non-CM rational points. The curve 48.144.8.lu.1 covers 24.36.1.fu.1 which is an elliptic curve of rank 0. It has exactly two rational points: one is a cusp and one is CM. The curve 48.144.9.zt.1 covers 24.36.1.fv.1 which is an elliptic curve of rank 0. It has exactly two rational points: one is a cusp and the other is CM. The curve 48.144.9.baj.1 has no $\mathbb{Q}_{5}$ points and thus no rational points. The curve 48.144.6.bk.2 covers 24.72.2.di.1 which is a genus 2, rank 2 bielliptic curve. It has model
$y^{2}=x^{6}+3x^{4}+3x^{2}+9.$
We know from Table 1 that it has exactly 8 rational points. All are CM points. The curve 48.144.9.bay.1 covers 24.36.1.ge.1 which, by Table LABEL:tab:coverings, has no non-CM rational points. The curve 48.144.10.lt.1 covers 24.12.1.bz.1 which has four rational points: two cusps and two CM points. The curve 48.144.7.mg.1 covers 48.24.1.f.2 which is an elliptic curve of rank 0. It has exactly two points: one is a cusp and one is CM. The curve 48.144.7.kt.1 covers 24.36.1.fr.1 which we already showed has no non-CM rational points. Similarly, the curve 48.144.7.mw.1 covers 24.36.1.gi.1 which we already showed has no non-CM rational points.
Although we have not determined the set of rational points of 48.72.3.a.1, we have now shown that none of its subgroups in our lattice could result in an adelic index outside $\mathcal{I}_{2}.$
We adopt a similar strategy for 48.72.3.a.2(=16.24.0.k.1 $\times\;K_{2}$). The children of 16.24.0.k.1 comprise the set {16.48.0.n.1, 16.48.0.h.1, 16.48.0.a.1, 16.48.1.bl.1, 16.48.0.m.2,
16.48.1.bg.1}. The labels of direct product of these with $K_{2}$ are {48.144.8.ca.1, 48.144.8.bd.2, 48.144.8.a.1, 48.144.7.pu.1, 48.144.8.bz.1, 48.144.7.pa.1}. We now discuss the rational points of these. The curves 48.144.8.ca.1, 48.144.8.bd.2, 48.144.8.a.1, 48.144.7.pu.1, 48.144.8.bz.1, and 48.144.7.pa.1 cover 12.36.1.bt.1, 12.36.1.bs.1, 12.36.2.a.1,
24.36.1.fp.1, 12.36.2.p.1, 24.36.1.fo.1, respectively. All of the latter curves have no non-CM rational points, so none of the former curves do either.
@@
We now discuss the computations of their rational points. The curve 48.144.9.zs.1 covers 24.36.1.fv.1 which is an elliptic curve of rank 0. It has exactly two rational points, one is a cusp and the other is CM. The curve 48.144.9.baz.1 covers 24.36.1.ge.1 which has no non-CM rational points. The curve 48.144.9.bai.1 has no $\mathbb{Q}_{5}$ points. The curve 48.144.6.bk.1 covers 24.72.2.di.1 which is a genus 2, rank 2 bielliptic curve. It has model
$y^{2}=x^{6}+3x^{4}+3x^{2}+9.$
By Table 1, it has exactly 8 rational points. All are CM points. The curve 48.144.7.hu.2 covers 12.36.1.bn.1 which has no non-CM rational points. The curve 48.144.8.lv.1 covers 24.36.1.fu.1 which is an elliptic curve of rank 0. It has exactly two rational points, one is a cusp and one is CM. The curve 48.144.7.mx.1 covers 24.36.1.gi.1 which we know already has no non-CM rational points. The curve 48.144.7.ks.1 covers 24.36.1.fr.1 which we already know has no non-CM rational points. The curve 48.144.7.mh.1 covers 24.72.2.gm.1 which is genus 2, rank 2 bielliptic. This curve has no rational points, which can be verified using a version of the Mordell–Weil sieve implemented by Bianchi and Padurariu designed to prove that the set of rational points for such curves is empty [simplesievecode]. The curve 48.144.10.ls.1 covers 24.24.1.dq.1 which is an elliptic curve of rank 0. It has exactly two rational points, both of which are CM.
To conclude, although we have not determined the set of rational points of 48.72.3.a.2, we have now shown that none of its subgroups in our lattice could result in an adelic index outside $\mathcal{I}_{2}.$
For 48.72.3.p.1 (=16.24.0.n.1 $\times$ $X_{\mathrm{ns}}^{+}(3)$), it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.a.1 that come from subgroups in our lattice. This is because, from Tables 11 and 12 of [MR4868206], if there is a non-CM $E/\mathbb{Q}$ whose $2$-adic image is a subgroup of 16.24.0.n.1, then there exists an $E^{\prime}/\mathbb{Q}$ that is 2-isogenous to $E$ whose $2$-adic image will be a subgroup of 16.24.0.k.2. From Lemma 8, admitting a 2-isogeny does not affect the $3$-adic image. So, if there is a non-CM $E/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.p.1, then there exists an $E^{\prime}/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.a.1. Further, from Corollary 4 of [MR4830946], we know that isogenous elliptic curves share the same adelic index. Lastly, by a similar argument, it suffices for the curve 48.72.3.p.2 to compute all the adelic indices of non-CM rational points of 48.72.3.a.2 that come from subgroups in our lattice.

### v2

@@ 8.2.2. Products with Infinitely Many Rational Points
24.36.0.cg.1 24.144.9.dga.1, 24.144.9.dgq.1, 24.144.9.eid.1, 24.144.9.ejb.1, 24.288.17.gil.1, 24.288.17.gim.1, 24.288.17.gim.2, 24.288.17.gin.1, 48.144.7.bfg.1, 48.144.7.bft.1, 48.144.7.ze.1, 48.144.7.zv.1, 48.144.9.bep.1, 48.144.9.bfa.1, 48.144.9.bjz.1, 48.144.9.bks.1, 48.144.11.sm.1, 48.144.11.sr.1, 48.288.21.gvf.1, 48.288.21.gvg.1, 48.288.21.gvg.2, 48.288.21.gvh.1,
24.36.0.cj.1 48.288.15.cva.1, 48.288.15.cvb.1, 48.288.15.cvz.1, 48.288.15.cwc.1, 48.288.15.cyo.1, 48.288.15.cyp.1, 48.288.15.cyv.1, 48.288.15.cyy.1
24.36.1.gl.1 48.288.15.cdy.1, 48.288.15.ceb.1, 48.288.17.fay.1, 48.288.17.faz.1, 48.288.19.hiq.1, 48.288.19.hit.1, 48.288.21.ezh.1, 48.288.21.ezi.1
48.72.0.c.1 48.144.3.bb.1, 48.144.3.r.1, 48.144.3.t.2, 48.144.3.z.1, 48.144.5.ee.2, 48.144.5.ef.1, 48.144.5.eg.1, 48.144.5.eh.2, 48.144.5.ei.2, 48.144.5.ej.1, 48.144.5.ek.1, 48.144.5.el.2, 48.144.5.em.1, 48.144.5.en.2, 48.144.5.eo.2, 48.144.5.ep.1, 48.144.5.eq.2, 48.144.5.er.1, 48.144.5.es.1, 48.144.5.et.2, 48.144.5.eu.1, 48.144.5.ev.2, 48.144.5.ew.2, 48.144.5.ex.1, 48.144.5.ey.2, 48.144.5.ez.1, 48.144.5.fa.1, 48.144.5.fb.2, 48.144.7.bas.1, 48.144.7.bgq.2, 48.144.11.ys.1, 48.144.11.yq.1
48.72.0.c.2 48.144.3.bb.2, 48.144.3.r.2, 48.144.3.t.1, 48.144.3.z.2, 48.144.5.ee.1, 48.144.5.ef.2, 48.144.5.eg.2, 48.144.5.eh.1, 48.144.5.ei.1, 48.144.5.ej.2, 48.144.5.ek.2, 48.144.5.el.1, 48.144.5.em.2, 48.144.5.en.1, 48.144.5.eo.1, 48.144.5.ep.2, 48.144.5.eq.1, 48.144.5.er.2, 48.144.5.es.2, 48.144.5.et.1, 48.144.5.eu.2, 48.144.5.ev.1, 48.144.5.ew.1, 48.144.5.ex.2, 48.144.5.ey.1, 48.144.5.ez.2, 48.144.5.fa.2, 48.144.5.fb.1, 48.144.7.bas.2, 48.144.7.bgq.1, 48.144.11.yq.2, 48.144.11.ys.2
48.72.0.d.1 48.144.3.bd.2, 48.144.3.bf.1, 48.144.5.fi.1, 48.144.5.fw.1, 48.144.5.fz.1, 48.144.5.gb.1, 48.144.5.gj.1, 48.144.5.gw.1, 48.144.5.gy.1, 48.144.5.hb.1
48.72.0.d.2 48.144.3.bd.1, 48.144.3.bf.2, 48.144.5.fm.1, 48.144.5.fx.1, 48.144.5.ga.1, 48.144.5.gd.1, 48.144.5.gf.1, 48.144.5.gu.1, 48.144.5.gx.1, 48.144.5.ha.1
These are mostly also amenable to techniques already discussed in this paper. We go through each in full detail in Section 8.2.m. We provide additional comments about curves with a more ad hoc approach here.
@@
The curve 48.144.3.bf.1 is a genus 3 rank 3 curve. Instead of computing its set of rational points explicitly, we check that it has commutator index 576 and compute its proper maximal subgroups that are fiber products of two respective factors, contain I-I, and have full determinant. There are 11 such groups whose labels are in the set {48.288.13.bhs.1, 48.288.13.bhr.1, 48.288.13.bib.1, 48.288.13.bhy.1, 48.288.21.hxr.1, 48.288.13.bft.1, 48.288.21.hxt.1, 48.288.13.bfu.1, 48.288.13.bga.1, 48.288.13.bgd.1, 48.288.21.hxl.2}. Each of these can be handled using methods already discussed, and this is also done in Section 8.2.m.
@@
The curve 48.144.3.bd.1 is a genus 3 rank 3 curve. Again, instead of computing its set of rational points explicitly, we check that it has commutator index 576 and compute its proper maximal subgroups that are fiber products of two respective factors that contain I-I and have full determinant. There are 11 such curves and their labels comprise the set {48.288.21.hxh.2, 48.288.13.bfj.1, 48.288.13.bfk.1, 48.288.13.bgs.1, 48.288.13.bgr.1, 48.288.13.bfa.1, 48.288.13.bgi.1, 48.288.13.bfd.1, 48.288.13.bgl.1, 48.288.21.hxf.1, 48.288.21.hxp.2}. We similarly handle these with discussed techniques on the GitHub file.
8.2.3. Rational points on 48.72.3.a.1,48.72.3.a.2,48.72.3.p.1, and 48.72.3.p.2
The curve 48.72.3.a.1 can be realized as 16.24.0.k.2 $\times\;K_{2}$. We divide our analysis into three parts. We first consider the direct product of children of 16.24.0.k.2 with $K_{2}$. Then, we consider the direct product of 16.24.0.k.2 with the proper subgroups of $K_{2}$ in this lattice. Lastly, we consider the proper maximal subgroups that are fiber products of the two respective factors that contain I-I, have full determinant, and could possibly contain a subgroup whose commutator index lies outside $\mathcal{I}_{2}$.
@@
There are six children of 16.24.0.k.2 with labels 16.48.0.n.1, 16.48.1.bq.1, 16.48.0.a.1, 16.48.0.h.2, 16.48.0.m.1, and 16.48.1.bv.1. The labels of the direct product of these with $K_{2}$ are 48.144.8.ca.1, 48.144.7.rm.1, 48.144.8.a.1, 48.144.8.bd.1, 48.144.8.bz.2, and 48.144.7.sg.1 respectively. Let us discuss their rational points. The curve 48.144.8.ca.1 covers 12.36.1.bt.1 which has rank 0. It has exactly 6 points and all are cusps or CM. The curve 48.144.7.rm.1 covers 24.72.2.jo.1 which is bielliptic and has rank 2. It has Weierstrass model
@@
For the second step, we consider the direct product of 16.24.0.k.2 with the proper subgroups of $K_{2}$ in this lattice; the only such subgroup is $K_{3}$. The direct product of 16.24.0.k.2 with $K_{3}$ has label 48.144.7.ik.1. This curve covers 12.36.1.bq.1 which we have already discussed: it has no non-CM rational points.
@@
In Section 8.2.m, we show that these curves all have no non-CM rational points. Although we have not determined the set of rational points of 48.72.3.a.1, we have now shown that none of its subgroups in our lattice could result in an adelic index outside $\mathcal{I}_{2}.$
We adopt a similar strategy as the previous subsection for 48.72.3.a.2, which can be realized as 16.24.0.k.1 $\times\;K_{2}$. The children of 16.24.0.k.1 comprise the set {16.48.0.n.1, 16.48.0.h.1, 16.48.0.a.1, 16.48.1.bl.1, 16.48.0.m.2, 16.48.1.bg.1}. The labels of direct product of these with $K_{2}$ are {48.144.8.ca.1, 48.144.8.bd.2, 48.144.8.a.1, 48.144.7.pu.1, 48.144.8.bz.1, 48.144.7.pa.1}. These cover 12.36.1.bt.1, 12.36.1.bs.1, 12.36.2.a.1, 24.36.1.fp.1, 12.36.2.p.1, and 24.36.1.fo.1, respectively. All of the latter curves have no non-CM rational points, so none of the former curves do either.
@@
We show in Section 8.2.m that these all have no non-CM rational points. These all use already-discussed methods, with the slight exception of the curve 48.144.7.mh.1. This covers 24.72.2.gm.1 which is genus 2, rank 2 bielliptic. This curve has no rational points, which can be verified using a version of the Mordell–Weil sieve implemented by Bianchi and Padurariu designed to prove that the set of rational points for such curves is empty [4]. To conclude, although we have not determined the set of rational points of 48.72.3.a.2, we have now shown that none of its subgroups in our lattice could result in an adelic index outside $\mathcal{I}_{2}.$
For 48.72.3.p.1 (=16.24.0.n.1 $\times$ $X_{\mathrm{ns}}^{+}(3)$), it suffices to compute all the adelic indices of non-CM rational points of 48.72.3.a.1 that come from subgroups in our lattice. This is because, from Tables 11 and 12 of [13], if there is a non-CM $E/\mathbb{Q}$ whose $2$-adic image is a subgroup of 16.24.0.n.1, then there exists an $E^{\prime}/\mathbb{Q}$ that is 2-isogenous to $E$ whose $2$-adic image will be a subgroup of 16.24.0.k.2. From Lemma 7, admitting a 2-isogeny does not affect the $3$-adic image. So, if there is a non-CM $E/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.p.1, then there exists an $E^{\prime}/\mathbb{Q}$ that corresponds to a rational point on 48.72.3.a.1. Further, from Corollary 4 of [21], we know that isogenous elliptic curves share the same adelic index. Lastly, by a similar argument, it suffices for the curve 48.72.3.p.2 to compute all the adelic indices of non-CM rational points of 48.72.3.a.2 that come from subgroups in our lattice.